OpenStudy (anonymous):

3. Tell which measure of central tendency best describes the data. Weights of books (oz): 12, 10, 9, 15, 16, 10 A) Mean B) Median C) Mode 4. Find the value of x such that the data set has the given mean. 99, 86, 76, 95, x; mean 91 A) x = 91 B) x = 87 C) x = 82 D) x = 99

OpenStudy (anonymous):

second one i can do

OpenStudy (anonymous):

the mean is 91 and you have 5 numbers (including \(x\)\) so you have to solve \[\frac{99+86+76+95+x}{5}=91\] add up top first, what do you get ?

OpenStudy (anonymous):

356x/5

OpenStudy (anonymous):

i hope you mean \[\frac{356+x}{5}\] right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so if \[\frac{356+x}{5}=91\] then \[365+x=91\times 5\]

OpenStudy (anonymous):

whatever that is then \[x=91\times 5-356\]

OpenStudy (anonymous):

i get \(99\) http://www.wolframalpha.com/input/?i=91*5-356

OpenStudy (anonymous):

btw pretty sure the first one is the mean, not the median or mode

OpenStudy (anonymous):

@satellite73 Thanks, could you help me with a few more?

OpenStudy (anonymous):

lol go ahead and post, i will look at them statistics is not really my forte though

OpenStudy (anonymous):

best to post in a new box to eliminate all the scrolling down also maybe you will get someone who knows what they are doing